# Markov-type state models to describe non-Markovian dynamics

### v1

@@ 3.2 The folding of HP35
Figure 2: Twelve-state MSM constructed from a $300\mu$s-long MD trajectory26 of the folding of HP35. (a) First two implied timescales obtained from the microstates (dashed black), the local equilibrium approximation (using the lag time $\tau=10\,$ns, black), the Hummer-Szabo projection (using $\tau=10\,$ns, green), the microstate-based result (blue), and the result from qMSM (using the kernel time $\tau_{\rm K}=10\,$ns, red). (b) Chapman-Kolmogorov test of the twelve macrostates, comparing the reference MD state trajectory to the various versions of the theory. Also shown are results of the hybrid MD/MSM calculation [Eq. (14)] with $t_{\rm max}=100\,$ns.
@@
To learn how the above findings for the toy model generalize to the case of all-atom MD data, we now consider the folding of villin headpiece (HP35) as a well-established model problem. As in previous work,28, 27 we use a $\sim 300\mu$s-long MD trajectory of the fast folding Lys24Nle/Lys29Nle mutant of HP35 at $T=360\,$K by Piana et al., 26 which shows about 30 folding events. Following the benchmark study of Nagel et al.,27 we employ the following MSM workflow. Firstly, we use MoSAIC correlation analysis29 to perform a feature selection resulting in 42 contact distances. Next, we eliminate high-frequency fluctuations of the distance trajectory, by employing a Gaussian low-pass filter28 with a standard deviation of $\sigma=2\,$ns. Employing principal component analysis on these contacts,30 we use the resulting first five components for subsequent robust density based clustering31 into 547 microstates. In a final step, we adopt the most probable path algorithm32 to dynamically lump the microstates into 12 metastable macrostates, using a lag time of 10 ns. The resulting dynamical model compares favorably to MSMs from alternative combinations of methods27 and explains the folding of HP35 as cooperative transition between folded and unfolded energy basins and several intermediate states. In this work, we employ the partitioning of the MD trajectory into micro- and macrostates, in order to calculate the various versions of the macrostate transition matrix introduced above.
We begin with the discussion of the first two implied timescales, which are depicted in Fig. 2a as a function of the lag time. As discussed in Ref. 27, the slowest timescale ($\sim 1.5\,\mu$s) corresponds to the overall folding process, while the next two timescales ($\sim 0.1\,\mu$s) account for conformational rearrangement in the native and the unfolded basins. Taking again the microstate timescales as a reference, we overall find the Hummer-Szabo results being in excellent agreement, followed by the results from the local equilibrium approximation, which virtually coincide with the qMSM and the microstate-based results.
As most important assessment, Fig. 2b shows the Chapman-Kolmogorov test of the twelve macrostates, which compare the macrostate projection of the MD trajectory (providing the reference) to the various versions of the theory. As a result of a careful selection of features and methods, the benchmark model27 represents a valuable MSM. Consequently, the deviations of the various methods are overall only minor, and already the simple local equilibrium approximation matches the correct population dynamics of most macrostates quite closely. Quite similar is the Hummer-Szabo projection, despite its superior modeling of the implied timescales above.
Recalling that the microstate-based calculation should coincide with the reference MD results in the case of a truly Markovian partitioning of the microstates, we learn that the microstates are a main reason for the deviations of the Chapman-Kolmogorov tests. Not relying on the quality of the microstates, the qMSM in fact represents an improvement of the microstate-based methods, particularly in the case of ill-defined states such as states 2 and 11. (Judging from the time evolution of the qMSM memory matrix ${\boldsymbol{K}}(t)$ shown in Fig. S1, we chose $\tau_{\rm K}=10$ns where all elements $K_{IJ}(t)$ are safely decayed.) Even better is the performance of a hybrid calculation [Eq. (14)], which uses until $t_{\rm max}=100\,$ns the exact transition matrix ${\boldsymbol{T}}_{\rm MD}(t)$, and for longer times the local approximation ${\boldsymbol{T}}_{\rm LE}(t=nt_{\rm max})$. Apart from the ill-defined states 2 and 11, the hybrid calculation matches the reference very accurately.
While a single, long MD trajectory is certainly perfect for building an MSM, in practice we often have many short trajectories (since they are readily computed in parallel). In principle, trajectories as short as the lag time $\tau$ and the kernel time $\tau_{\rm K}$ should be already sufficient to construct an MSM and a qMSM, respectively. However, the validity of this presumption has been rarely assessed due to the lack of long reference trajectories. As we used $\tau=\tau_{\rm K}=10\,$ns in the calculations above, and we need some extra time for the time averages to calculate transition probabilities, we decided to split the $300\mu$s trajectory in $10~{}000\times 30\,$ns (150 frames) long pieces, such that we can average over 100 frames. Note that 30 ns is quite short considering the overall folding time of $\sim 2\,\mu$s of HP35. As the MSM workflow explained above employs no time information up to the density based clustering, 33 we can use the same microstates. On the other hand, the dynamical lumping of these states via the most probable path algorithm32 requires calculations based on the transition matrix, which differs in the case of one long and of many short trajectories. Hence, the resulting macrostates differ minor from the ones obtained from the single $300\mu$s trajectory, see Fig. S2. Lastly, we again calculated the macrostate transition matrices for the various theoretical formulations.
Figure 3 shows the implied timescales and Chapman-Kolmogorov tests of the resulting MSMs. Overall the short-trajectories results look very similar to the results from the single long trajectory (Fig. 2), thus confirming the promise of MSMs to describe long-time dynamics from short trajectories. However, it should be kept in mind that these short trajectories used adequately chosen initial conditions, because they were taken from a long trajectory that exhibits numerous folding and unfolding transition. This is less obvious without such a reactive trajectory, although various strategies have been proposed to obtain good initial conditions along the reaction paths of a considered process.34, 35, 36
@@
Figure 3: As in Fig. 2, except that the original $300\mu$s-long trajectory of HP35 was split in $10~{}000\times 30\,$ns trajectories. The hybrid MD/MSM method is not shown here, as $t_{\textit{max}}$ would be too short in this case, due to poor statistics.
@@ 4 Concluding remarks
We have outlined the theoretical basis of various approaches to estimate the macrostate transition matrix of an MSM, and discussed their performance using a one-dimensional toy model and an all-atom folding trajectory of HP35. The most commonly used method, the local equilibrium approximation [Eq. (8)], simply counts the transitions between the macrostates within a certain lag time, which assumes Markovian data with a timescale separation between fast intrastate fluctuations and slow interstate transitions. As shown for the toy model in Fig. 1, which allows to tune the ratio between these timescales, the quality of the resulting implied timescale and the Chapman-Kolmogorov test readily deteriorate, when this separation is not fulfilled.
Assuming that Markovianity is at least given at the microstate level, the knowledge of the microstate dynamics can be exploited to construct improved macrostate dynamics. For one, this is used in the Laplace-transform based method by Hummer and Szabo, 13 which was shown to achieve significantly improved Markovianity in all considered cases. However, we can also directly employ a microstate-to-macrostate projection to construct a macrostate transition matrix [Eq. (9)], which by design yields correct macrostate population dynamics. Moreover, this microstate-based approach may provide state-to-state waiting times and transition pathways, simply by performing microstate Monte Carlo Markov chain simulations with subsequent projection on macrostates. If the formulation does not achieve perfect Chapman-Kolmogorov tests, we know that this must be caused by non-Markovian microstate dynamics.
@@
In practice, an MSM is often used to interpret given MD data in terms of the time evolution of the populations $P_{I}(t)$ of metastable conformational states. For the length of the MD trajectories $t_{\rm max}$ (actually some hundreds MD frames less to facilitate a sufficient time average), we can simply project the MD data on these states, without invoking a dynamical approximation such as the assumption of Markovianity. In the ‘hybrid MD/MSM’ method [Eq. (14)], we exploit this idea to obtain the system’s short time evolution with good time resolution, and subsequently use some formulation (including local equilibrium, Hummer-Szabo, or microstate-based) to construct a MSM using the long lag time $t_{\rm max}$. Combining the best of two worlds, this strategy typically gives prime results.
@@ Acknowledgments
The authors thank Georg Diez, Daniel Nagel and Tim Uttenweiler for helpful comments and discussions, as well as D. E. Shaw Research for sharing their trajectories of HP35. This work has been supported by the Deutsche Forschungsgemeinschaft (DFG) within the framework of the Research Unit FOR 5099 ”Reducing complexity of nonequilibrium” (project No. 431945604), the High Performance and Cloud Computing Group at the Zentrum für Datenverarbeitung of the University of Tübingen, the state of Baden-Württemberg through bwHPC and the DFG through grant no INST 37/935-1 FUGG (RV bw16I016), the Black Forest Grid Initiative, and the Freiburg Institute for Advanced Studies (FRIAS) of the Albert-Ludwigs-University Freiburg.

### v2

@@ 3.2 The folding of HP35
Figure 2: Twelve-state MSM constructed from a $300\mu$s-long MD trajectory30 of the folding of HP35. (a) First two implied timescales obtained from the microstates (dashed black, reference), the local equilibrium approximation (using the lag time $\tau=10\,$ns, black), the Hummer-Szabo projection (using $\tau=10\,$ns, green), the microstate-based result (blue), and the result from qMSM (using the kernel time $\tau_{\rm K}=10\,$ns, red). (b) Chapman-Kolmogorov test of the twelve macrostates, comparing the reference MD state trajectory to the various versions of the theory. Also shown are results of the hybrid MD/MSM calculation [Eq. (14)] with $t_{\rm max}=100\,$ns.
@@
To learn how the above findings for the toy model generalize to the case of all-atom MD data, we now consider the folding of villin headpiece (HP35) as a well-established model problem. As in previous work,35, 33 we use a $\sim 300\mu$s-long MD trajectory of the fast folding Lys24Nle/Lys29Nle mutant of HP35 at $T=360\,$K by Piana et al., 30 which shows about 30 folding events. Following the benchmark study of Nagel et al.,33 we employ the following MSM workflow. Firstly, we use MoSAIC correlation analysis36 to perform a feature selection resulting in 42 contact distances. Next, we eliminate high-frequency fluctuations of the distance trajectory, by employing a Gaussian low-pass filter35 with a standard deviation of $\sigma=2\,$ns. Employing principal component analysis on these contacts,37 we use the resulting first five components for subsequent robust density based clustering38 into 547 microstates. In a final step, we adopt the most probable path algorithm39 to dynamically lump the microstates into 12 metastable macrostates, using a lag time of 10 ns. The resulting dynamical model compares favorably to MSMs from alternative combinations of methods33 and explains the folding of HP35 as cooperative transition between folded and unfolded energy basins and several intermediate states. In this work, we employ the partitioning of the MD trajectory into micro- and macrostates, in order to calculate the various versions of the macrostate transition matrix introduced above.
@@
Figure 3: As in Fig. 2, except that the original $300\mu$s-long trajectory of HP35 was split in $10~{}000\times 30\,$ns trajectories. The hybrid MD/MSM method is not shown, as for $t_{\textit{max}}=10\,$ns it coincides with the local equilibrium approximation.
We begin with the discussion of the first two implied timescales, which are depicted in Fig. 2a as a function of the lag time. As discussed in Ref.  33, the slowest timescale ($\sim 1.5\,\mu$s) corresponds to the overall folding process, while the next two timescales ($\sim 0.1\,\mu$s) account for conformational rearrangement in the native and the unfolded basins. Taking again the microstate timescales as a reference, we overall find the Hummer-Szabo results being in excellent agreement, followed by the results from the local equilibrium approximation, which virtually coincide with the qMSM and the microstate-based results.
As most important assessment, Fig. 2b shows the Chapman-Kolmogorov test of the twelve macrostates, which compare the macrostate projection of the MD trajectory (providing the reference) to the various versions of the theory. As a result of a careful selection of features and methods, the benchmark model33 represents a valuable MSM. Consequently, the deviations of the various methods are overall only minor, and already the simple local equilibrium approximation matches the correct population dynamics of most macrostates quite closely. Quite similar is the Hummer-Szabo projection, despite its superior modeling of the implied timescales above.
Recalling that the microstate-based calculation should coincide with the reference MD results in the case of a truly Markovian partitioning of the microstates, we learn that the microstates are a main reason for the deviations of the Chapman-Kolmogorov tests. Not relying on the quality of the microstates, the qMSM in fact represents an improvement of the microstate-based methods, particularly in the case of ill-defined states such as states 2 and 11. (Judging from the time evolution of the qMSM memory matrix ${\boldsymbol{K}}(t)$ shown in Fig. S2, we chose $\tau_{\rm K}=10\,$ns where, apart from residual fluctuations around zero, all elements $K_{IJ}(t)$ are safely decayed.) Even better is the performance of a hybrid calculation [Eq. (14)], which uses until $t_{\rm max}=100\,$ns the exact transition matrix ${\boldsymbol{T}}_{\rm MD}(t)$, and for longer times the local approximation ${\boldsymbol{T}}_{\rm LE}(t=mt_{\rm max})$. Apart from the ill-defined states 2 and 11, the hybrid calculation matches the reference very accurately.
While a single, long MD trajectory is certainly perfect for building an MSM, in practice we often have many short trajectories (since they are readily computed in parallel). In principle, trajectories as short as the lag time $\tau$ and the kernel time $\tau_{\rm K}$ should be already sufficient to construct an MSM and a qMSM, respectively. However, the validity of this presumption has been rarely assessed due to the lack of long reference trajectories. As we used $\tau=\tau_{\rm K}=10\,$ns in the calculations above, and we need some extra time for the time averages to calculate transition probabilities, we decided to split the $300\mu$s trajectory in $10000\times 30\,$ns (150 frames) long pieces, such that we can average over 100 frames. Note that 30 ns is quite short considering the overall folding time of $\sim 2\,\mu$s of HP35. As the MSM workflow explained above employs no time information up to the density based clustering included, 40 we can use the same microstates. On the other hand, the dynamical lumping of these states via the most probable path algorithm39 requires calculations based on the transition matrix, which differs in the case of one long and of many short trajectories. Hence, the resulting macrostates differ minor from the ones obtained from the single $300\mu$s trajectory, see Fig. S3. Lastly, we again calculated the macrostate transition matrices for the various theoretical formulations.
Figure 3 shows the implied timescales and Chapman-Kolmogorov tests of the resulting MSMs. Overall the short-trajectories results look very similar to the results from the single long trajectory (Fig. 2), thus confirming the promise of MSMs to describe long-time dynamics from short trajectories. However, it should be kept in mind that these short trajectories used adequately chosen initial conditions, because they were taken from a long trajectory that exhibits numerous folding and unfolding transition. This is less obvious without such a reactive trajectory, although various strategies have been proposed to obtain good initial conditions along the reaction paths of a considered process.41, 42, 43
@@ 4 Concluding remarks
We have outlined the theoretical basis of various approaches to estimate the macrostate transition matrix of an MSM, from the standard local equilibrium approximation to a sophisticated generalized master equation approach. Employing a one-dimensional toy model and an all-atom folding trajectory of HP35, we have provided a comprehensive comparison of the various formulations, and introduced two new methods, the microstate-based approach in Eq. (9) and the hybrid MD/MSM ansatz in Eq. (14).
The most commonly used method, the local equilibrium approximation [Eq. (8)], simply counts the transitions between the macrostates within a certain lag time, which assumes Markovian data with a timescale separation between fast intrastate fluctuations and slow interstate transitions. As shown for the toy model in Fig. 1, which allows to tune the ratio between these timescales, the quality of the resulting implied timescale and the Chapman-Kolmogorov test readily deteriorate, when this separation is not fulfilled.
Assuming that Markovianity is at least given at the microstate level, the knowledge of the microstate dynamics can be exploited to construct improved macrostate dynamics. For one, this is used in the Laplace transform-based method by Hummer and Szabo, 13 which was shown to achieve significantly improved Markovianity in all considered cases. However, we can also directly employ a microstate-to-macrostate projection to construct a macrostate transition matrix ${\boldsymbol{T}}_{\rm Mic}(t)$ [Eq. (9)], which by design yields correct macrostate population dynamics. As an important consequence, this means that residual non-Markovian behavior of the resulting MSM must be caused by suboptimal microstates (and not by the lumping of micro- into macrostates). By performing microstate Monte Carlo Markov chain simulations with subsequent projection on macrostates, this new and promising microstate-based approach may also provide state-to-state waiting times and transition pathways.
@@
In practice, an MSM is often used to interpret given MD data in terms of the time evolution of the populations $P_{I}(t)$ of metastable conformational states. For the length of the MD trajectories $t_{\rm max}$ (actually some hundreds MD frames less to facilitate a sufficient time average), we can simply project the MD data on these states, without invoking a dynamical approximation such as the assumption of Markovianity. In the ‘hybrid MD/MSM’ method [Eq. (14)], we exploit this idea to obtain the system’s short time evolution with good time resolution, and subsequently use some formulation (including local equilibrium, Hummer-Szabo, or microstate-based) to construct a MSM using the long lag time $t_{\rm max}$. Given sufficiently long trajectories, this strategy of combining the best of two worlds is simple and promises prime results. While the definition of the hybrid MD/MSM method and its virtues seem obvious, it has to the best of our knowledge not yet been fully appreciated.
@@ Acknowledgments
The authors thank Georg Diez, Daniel Nagel and Tim Uttenweiler for helpful comments and discussions, as well as D. E. Shaw Research for sharing their trajectories of HP35. This work has been supported by the Deutsche Forschungsgemeinschaft (DFG) within the framework of the Research Unit FOR 5099 ”Reducing complexity of nonequilibrium” (project No. 431945604), the High Performance and Cloud Computing Group at the Zentrum für Datenverarbeitung of the University of Tübingen and the Rechenzentrum of the University of Freiburg, the state of Baden-Württemberg through bwHPC and the DFG through Grant Nos. INST 37/935-1 FUGG (RV bw16I016) and INST 39/963-1 FUGG (RV bw18A004), the Black Forest Grid Initiative, and the Freiburg Institute for Advanced Studies (FRIAS) of the Albert-Ludwigs-University Freiburg.